OpenStudy (anonymous):

An 18 ouncw box of cereal costs 5.40 . How much would you expect to pay if you bought the 24 ounce box?

OpenStudy (anonymous):

you have many ways to do this but all involve ratios

OpenStudy (anonymous):

i know HELP

OpenStudy (anonymous):

\[\frac{5.40}{18}=\frac{x}{24}\]

OpenStudy (anonymous):

that would be a start. \[\frac{price}{ounce}=\frac{price}{ounce}\]

OpenStudy (anonymous):

but how much money would it cost

OpenStudy (anonymous):

The ratio of the box masses should be the same as the ratio of the costs (unless you get a bulk discount), i.e. \[\frac{24}{18} = \frac{x}{5.40}\]

OpenStudy (anonymous):

so answer is \[x=\frac{24\times 5.4}{18}\] i get $7.20

OpenStudy (anonymous):

UR RIGHT

OpenStudy (anonymous):

give me a medal for trying

OpenStudy (anonymous):

see i told you several ways to do it. notice that solving the way dalvoron wrote it also involves computing \[x=\frac{24\times 5.4}{18}\]

OpenStudy (anonymous):

$7.20

OpenStudy (zarkon):

I would expect to pay less than $7.20 Shouldn't you get a discount when you buy more :)